OpenStudy (anonymous):

What is 64i in radical form?

OpenStudy (perl):

im not sure what you mean by radical form

OpenStudy (anonymous):

like how 4i√2 would be √-32.

OpenStudy (anonymous):

\[ 64i = \sqrt{x} \implies x = 64^2 i^2 = -64^2. \]Can you finish now?